OpenStudy (anonymous):

if z is acomplex number that satisfies the equation z^4+z^3+2z^2+z+1=0,then find the value of the conjugate of z

OpenStudy (perl):

so assuming z is a solution, we want z* the conjugate

OpenStudy (perl):

if f (z ) = 0, then f ( z* ) = 0

OpenStudy (perl):

we might be able to multiply ( z^4+z^3+2z^2+z+1) * ( z'^4+z'^3+2z'^2+z'+1) = 0 * 0

OpenStudy (perl):

i will put z = a + bi, and z ' = a - bi , into wolfram, and see what it says

OpenStudy (perl):

z= i, and z = -i solve the equation

OpenStudy (anonymous):

yeah when i substitute z as i ,the equation is right.....but how do you figure it up?

OpenStudy (perl):

but i think we should use properties of complex numbers, do you have any notes on them

OpenStudy (anonymous):

i don't

OpenStudy (zarkon):

\[z^4+z^3+2z^2+z+1=z^4+z^3+z^2+z^2+z+1\] \[(z^4+z^3+z^2)+(z^2+z+1)=z^2(z^2+z+1)+(z^2+z+1)\] \[=(z^2+1) (z^2+z+1)\]